# Think Before Recommend: Unleashing the Latent Reasoning Power for Sequential Recommendation

### v2

@@ 3.1.2. Extended Inference-Time Reasoning
To differentiate between item encoding outputs and reasoning outputs, we denote the hidden states of the model’s final layer from position to as , where represents the reasoning hidden state at the -th step. To obtain the user representation, a straightforward approach is to follow the traditional paradigm i.e., use the last reasoning output as . Then, we calculate the predicted probability for the user as and use cross-entropy loss as the recommendation objective function:
@@ 3.2.1. Multi-Step Reasoning Supervision
Specifically, we treat the reasoning hidden states from multiple steps as multi-vector user representations and apply separate recommendation losses to each in order to enhance process guidance:
where stands for the prediction probability of the ground-truth item based on the -th reasoning output, i.e.:
@@ 3.3.1. Progressive Temperature Annealing (PTA)
Next, we continue to use Eq. (6) to inject the process supervisor into the reasoning process. With this lean annealing strategy, the model is encouraged to explore a broader solution space in the early reasoning stage, preventing it from getting stuck in local optima. Then, as the reasoning process progresses, the value of is gradually reduced to narrow the search space, guiding the model towards the global optimum. Thus, the proposed PTA can more effectively approximate the user’s true preference distribution.
@@ 4. Experiments
Dataset | Yelp | Video & Games | Software | CDs & Vinyl | Baby & Products |
#Users | 13,083 | 89,021 | 30,049 | 35,238 | 140,292 |
#Items | 10,697 | 22,933 | 16,705 | 87,969 | 30,689 |
#Avg. Inter. / User | 33.92 | 5.96 | 5.59 | 14.59 | 5.57 |
#Avg. Inter. / Item | 41.49 | 23.15 | 10.06 | 5.84 | 25.44 |
#Avg. Inter. | 443,807 | 530,989 | 168,029 | 513,991 | 780,809 |
@@ 4.1.2. Evaluation Metrics
Dataset | Method | SASRec | BERT4Rec | |||||||||
N@10 | N@20 | R@10 | R@20 | Avg. | N@10 | N@20 | R@10 | R@20 | Avg. | |||
Yelp | Base | 0.0347 | 0.0452 | 0.0626 | 0.1047 | - | 0.0364 | 0.046 | 0.0653 | 0.1038 | - | |
+ERL (Improv.) | 0.0383 (10.37%) | 0.0474 (4.87%) | 0.0691 (10.38%) | 0.1056 (0.86%) | 6.62% | 0.0371 (1.92%) | 0.0476 (3.48%) | 0.0661 (1.23%) | 0.1077 (3.76%) | 2.60% | ||
+PRL (Improv.) | 0.0388 (11.82%) | 0.0493 (9.07%) | 0.073 (16.61%) | 0.1149 (9.74%) | 11.81% | 0.0377 (3.57%) | 0.0487 (5.87%) | 0.0708 (8.42%) | 0.1149 (10.69%) | 7.14% | ||
Video & Games | Base | 0.0284 | 0.0353 | 0.0542 | 0.0816 | - | 0.0289 | 0.0355 | 0.0548 | 0.0810 | - | |
+ERL (Improv.) | 0.0301 (5.99%) | 0.0385 (9.07%) | 0.0581 (7.20%) | 0.0915 (12.13%) | 8.59% | 0.0311 (7.61%) | 0.0375 (5.63%) | 0.0578 (5.47%) | 0.0832 (2.72%) | 5.36% | ||
+PRL (Improv.) | 0.0299 (5.28%) | 0.0379 (7.37%) | 0.0572 (5.54%) | 0.0890 (9.07%) | 6.81% | 0.0306 (5.88%) | 0.0380 (7.04%) | 0.0584 (6.57%) | 0.0879 (8.52%) | 7.00% | ||
Software | Base | 0.0696 | 0.0895 | 0.1468 | 0.2264 | - | 0.0710 | 0.0893 | 0.1530 | 0.2258 | - | |
+ERL (Improv.) | 0.0743 (6.75%) | 0.0935 (4.47%) | 0.1456 (0.82%) | 0.2224 (1.77%) | 2.16% | 0.0769 (8.31%) | 0.0964 (7.95%) | 0.1554 (1.57%) | 0.2328 (3.10%) | 5.23% | ||
+PRL (Improv.) | 0.0739 (6.18%) | 0.0949 (6.03%) | 0.1488 (1.36%) | 0.2324 (2.65%) | 4.06% | 0.0762 (7.32%) | 0.0976 (9.29%) | 0.1500 (1.96%) | 0.2350 (4.07%) | 4.68% | ||
CDs & Vinyl | Base | 0.0148 | 0.0174 | 0.0317 | 0.0419 | - | 0.0149 | 0.0185 | 0.0326 | 0.0468 | - | |
+ERL (Improv.) | 0.0182 (22.97%) | 0.0212 (21.84%) | 0.0363 (14.51%) | 0.0482 (15.04%) | 18.59% | 0.0165 (10.74%) | 0.0208 (12.43%) | 0.0354 (8.59%) | 0.0524 (11.97%) | 10.93% | ||
+PRL (Improv.) | 0.0155 (4.73%) | 0.0195 (12.07%) | 0.0315 (0.63%) | 0.0470 (12.17%) | 7.08% | 0.0162 (8.72%) | 0.0202 (9.19%) | 0.0334 (2.45%) | 0.0496 (5.98%) | 6.59% | ||
Baby & Products | Base | 0.0112 | 0.0157 | 0.0260 | 0.0437 | - | 0.0109 | 0.0154 | 0.0257 | 0.0439 | - | |
+ERL (Improv.) | 0.0116 (3.57%) | 0.0164 (4.46%) | 0.0228 (12.31%) | 0.0418 (4.35%) | 2.16% | 0.0148 (35.78%) | 0.0195 (26.62%) | 0.0293 (9.57%) | 0.0481 (14.01%) | 21.49% | ||
+PRL (Improv.) | 0.0135 (20.54%) | 0.0178 (13.38%) | 0.0281 (8.08%) | 0.0451 (3.20%) | 11.30% | 0.0140 (28.44%) | 0.0185 (20.13%) | 0.0291 (6.15%) | 0.0466 (13.23%) | 16.99% |
Dataset | Method | UniSRec | MoRec | |||||||||
N@10 | N@20 | R@10 | R@20 | Avg. | N@10 | N@20 | R@10 | R@20 | Avg. | |||
Yelp | Base | 0.0380 | 0.0495 | 0.0737 | 0.1195 | - | 0.0391 | 0.0516 | 0.0757 | 0.1258 | - | |
+ERL (Improv.) | 0.0406 (6.84%) | 0.0521 (5.25%) | 0.0770 (4.48%) | 0.1227 (2.68%) | 4.81% | 0.0417 (6.65%) | 0.0531 (2.91%) | 0.0832 (9.91%) | 0.1283 (1.99%) | 5.36% | ||
+PRL (Improv.) | 0.0413 (8.68%) | 0.0529 (6.87%) | 0.0788 (6.92%) | 0.1253 (4.85%) | 6.83% | 0.0410 (4.86%) | 0.0532 (3.10%) | 0.0804 (6.21%) | 0.1289 (2.46%) | 4.16% | ||
Video & Games | Base | 0.0328 | 0.0421 | 0.0683 | 0.1054 | - | 0.0350 | 0.0438 | 0.0716 | 0.1065 | - | |
+ERL (Improv.) | 0.0364 (10.98%) | 0.0440 (4.51%) | 0.0711 (4.10%) | 0.1015 (3.70%) | 3.97% | 0.0392 (12.00%) | 0.0485 (10.73%) | 0.0744 (3.91%) | 0.1112 (4.41%) | 7.76% | ||
+PRL (Improv.) | 0.0352 (7.32%) | 0.0433 (2.85%) | 0.0658 (3.66%) | 0.0982 (6.83%) | 0.08% | 0.0371 (6.00%) | 0.0462 (5.48%) | 0.0708 (1.12%) | 0.1067 (0.19%) | 2.64% | ||
Software | Base | 0.0820 | 0.1041 | 0.1643 | 0.2522 | - | 0.0846 | 0.1050 | 0.1697 | 0.2510 | - | |
+ERL (Improv.) | 0.0851 (3.78%) | 0.1075 (3.27%) | 0.1669 (1.58%) | 0.2556 (1.35%) | 2.49% | 0.0881 (4.14%) | 0.1071 (2.00%) | 0.1711 (0.82%) | 0.2466 (1.75%) | 1.30% | ||
+PRL (Improv.) | 0.0869 (5.98%) | 0.1076 (3.36%) | 0.1687 (2.68%) | 0.2518 (0.16%) | 2.96% | 0.0917 (8.39%) | 0.1120 (6.67%) | 0.1723 (1.53%) | 0.2532 (0.88%) | 4.37% | ||
CDs & Vinyl | Base | 0.0150 | 0.0208 | 0.0298 | 0.0527 | - | 0.0186 | 0.0235 | 0.0405 | 0.0604 | - | |
+ERL (Improv.) | 0.0208 (38.67%) | 0.0259 (24.52%) | 0.0428 (43.62%) | 0.0629 (19.35%) | 31.54% | 0.0199 (6.99%) | 0.0248 (5.53%) | 0.0417 (2.96%) | 0.0609 (0.83%) | 4.08% | ||
+PRL (Improv.) | 0.0191 (27.33%) | 0.0253 (21.63%) | 0.0394 (32.21%) | 0.0640 (21.44%) | 25.66% | 0.0198 (6.45%) | 0.0249 (5.96%) | 0.0417 (2.96%) | 0.0618 (2.32%) | 4.42% | ||
Baby & Products | Base | 0.0152 | 0.0199 | 0.0315 | 0.0501 | - | 0.0176 | 0.0231 | 0.0371 | 0.0588 | - | |
+ERL (Improv.) | 0.0183 (20.39%) | 0.0239 (20.10%) | 0.0367 (16.51%) | 0.0589 (17.56%) | 18.64% | 0.0184 (4.55%) | 0.0242 (4.76%) | 0.0373 (0.54%) | 0.0602 (2.38%) | 3.06% | ||
+PRL (Improv.) | 0.0182 (19.74%) | 0.0236 (18.59%) | 0.0359 (13.97%) | 0.0575 (14.77%) | 16.77% | 0.0189 (7.39%) | 0.0247 (6.93%) | 0.0376 (1.35%) | 0.0611 (3.91%) | 4.89% |
@@ 4.1.4. Implementation Details
We conduct all experiments on 8 NVIDIA A100 GPUs. To ensure a fair comparison, we set the embedding size and batch size for all methods to 256 and 2048, respectively. We optimize all models using the Adam (Kingma and Ba, 2014) optimizer with a learning rate of 0.001 and follow previous work (Sun et al., 2019) by adopting GeLU as the activation function. Following the existing works (Xie et al., 2022; Chen et al., 2022), we truncate user sequences to a maximum length of 50 across all datasets. Since our framework is model-agnostic, it can be seamlessly integrated into various sequential recommendation models. In particular, for BERT4Rec’s bidirectional Transformer, we employ a Prefix Masking strategy, where the item sequence part utilizes bidirectional attention, while the reasoning adopts unidirectional attention. Early stopping is triggered if the metrics on the validation set do not improve over 10 consecutive epochs. For item-based methods, we follow previous work (Liu et al., 2024) by using LLaMA-3.1-8B (Grattafiori et al., 2024) to encode item textual features. In particular, we apply Principle Component Analysis (PCA) to the averaged hidden states from the last layer, preserving core features and distilling 768-dimensional model representations. For ERL method, we search for the KL regularization hyperparameter within . For PRL method, we set the noise strength and tune the base temperature and temperature decay rate over the ranges and , respectively. Our code will be available at https://github.com/TangJiakai/ReaRec.
@@ 5.1. Sequential Recommendation
Sequential recommendation, as one of the core tasks within the field of Recommender Systems (RS), aims to predict the next item with which a user may interact by modeling behavior patterns and evolving user interests from user-item interaction sequences (Boka et al., 2024; Wang et al., 2019; Fang et al., 2020). Early studies focused on item-to-item collaborative transitions, typically employing Markov Chain-based matrix factorization methods (He and McAuley, 2016; Rendle et al., 2010) to achieve next-item predictions. With the advent of deep learning, researchers began exploring various sequential modeling architectures, such as Recurrent Neural Networks (RNNs) (Hidasi et al., 2015, 2016; Quadrana et al., 2017), Convolutional Neural Networks (CNNs) (Tang and Wang, 2018; Yuan et al., 2019; Yan et al., 2019), and Transformer (Kang and McAuley, 2018; Chen et al., 2019; Yang et al., 2022), to enhance modeling capabilities. Specifically, GRU4Rec (Hidasi et al., 2015) first introduced the GRU networks for session-based recommendation. Caser (Tang and Wang, 2018) applied convolutional operations by treating the item sequence embedding matrix as an “image” to extract multi-level interaction features. With the emergence of the Transformer architecture, numerous studies in sequential recommendation shifted toward self-attention-based models. For instance, SASRec (Kang and McAuley, 2018), a classic sequential recommendation baseline, incorporated self-attention to automatically learn the importance weights of historical items. BERT4Rec (Sun et al., 2019) further advanced this by adopting bidirectional encoding to capture more contextual dependencies from both directions of the sequence. Furthermore, to address data sparsity and cold-start issues, another line of research focused on leveraging common item attributes (such as text and images) to learn universal item and sequence representations (Hou et al., 2022; Yuan et al., 2023; Ding et al., 2021; Hou et al., 2023). For example, UniSRec (Hou et al., 2022) utilized multi-domain recommendation data to learn transferable sequential models. However, existing methods remain constrained by reasoning-free forward computation. In this paper, we further explore the potential reasoning capabilities of sequential recommenders by extending the multi-step computational depth at inference time.
5.2. Inference-time Reasoning
@@
As the scaling law of large language models (LLMs) at the training stage has gradually reached its bottleneck (Chen et al., 2025; Snell et al., 2024; Kumar et al., 2025), researchers have shifted their focus toward inference-time scaling. Notably, OpenAI’s O1 series (Jaech et al., 2024), Qwen’s QwQ series (Team, 2024), and DeepSeek’s R1 series (Guo et al., 2025) have emerged as key milestone works, marking the transition from conversational AI to reasoning-intensive AI and opening promising pathways toward achieving Artificial General Intelligence (AGI). These works leverage emerging long Chain-of-Thought (CoT) mechanisms to reveal excellent test-time scaling phenomena—where increased computational power (via generating more tokens) during inference substantially improves the model’s problem-solving abilities (Wei et al., 2022; Besta et al., 2024; Yao et al., 2023; Yang et al., 2024; Wang et al., 2022; Chu et al., 2023; Zhang et al., 2025). Extended reasoning space enables depth-scalable exploration, manifesting self-reflection capabilities (e.g., “Aha Moments”) that surpass traditional short-chain reasoning limitations (Chen et al., 2025; Ji et al., 2025). Compared with prior methods constrained by token-by-token generation in the discrete language spaces, which restricts the model’s expressive ability, another research direction focuses on implicit chain of thought reasoning in latent spaces, achieving both efficiency and performance gains in large language models (Hao et al., 2024; Xu et al., 2025a; Geiping et al., 2025; Chen et al., 2024) and multimodal foundation models (Shen et al., 2025; He et al., 2024). For example, Coconut (Hao et al., 2024) introduces continuous thinking in the latent reasoning space of LLMs, while Heima (Shen et al., 2025) compresses the entire multimodal CoT process into a single high-level thinking token (i.e., ¡CoT¿) to eliminate redundant intermediate token generation. Inspired by this think-before-action paradigm, we pioneer the exploration of implicit reasoning-enhanced sequential recommendation framework in this paper, proposing two lightweight reasoning learning methods to push the performance ceiling of sequential recommenders.

### v3

@@ 3.1.2. Extended Inference-Time Reasoning
To differentiate between item encoding outputs and reasoning outputs, we denote the hidden states of the model’s final layer from position to as , where represents the reasoning hidden state at the -th step. To obtain the user representation, a straightforward approach is to follow the traditional paradigm i.e., use the last reasoning output as . Then, we calculate the predicted probability for the user as
and use cross-entropy loss as the recommendation objective:
@@ 3.2.1. Multi-Step Reasoning Supervision
Specifically, we treat the reasoning hidden states from multiple steps as multi-vector user representations and apply cross-entropy loss (cf. Eq. (6)) to the ensembled sequence representation to enhance process guidance. Therefore, instead of using only the reasoning state from the last step, ERL employs an average pooling layer to aggregate reasoning hidden states across all steps to obtain the final user representation as . The output distribution is then computed following Eq. (5).
@@ 3.2.2. KL Divergence Regularization
where represents the logit output of the -th reasoning step.
@@ 3.3.1. Progressive Temperature Annealing (PTA)
In contrast to ensemble reasoning learning method, we apply separate recommendation losses to each reasoning hidden state to inject process supervision into the reasoning process, as follows:
where represents the logit corresponding to the item. With this lean annealing strategy, the model is encouraged to explore a broader solution space in the early reasoning stage, preventing it from getting stuck in local optima. Then, as the reasoning process progresses, the value of is gradually reduced to narrow the search space, guiding the model towards the global optimum. Thus, the proposed PTA can more effectively approximate the user’s true preference distribution.
@@ 4. Experiments
| Dataset | Yelp |
Software |
| #Users | 13,083 | 89,021 | 30,049 | 35,238 | 140,292 | ||||||
| #Items | 10,697 | 22,933 | 16,705 | 87,969 | 30,689 | ||||||
| #Avg. Inter. / User | 33.92 | 5.96 | 5.59 | 14.59 | 5.57 | ||||||
| #Avg. Inter. / Item | 41.49 | 23.15 | 10.06 | 5.84 | 25.44 | ||||||
| #Avg. Inter. | 443,807 | 530,989 | 168,029 | 513,991 | 780,809 | ||||||
@@ 4.1.2. Evaluation Metrics
| Dataset | Method | SASRec | BERT4Rec | ||||||||||||||||||||||||||
| N@10 | N@20 | R@10 | R@20 | Avg. | N@10 | N@20 | R@10 | R@20 | Avg. | ||||||||||||||||||||
| Yelp | Base | 0.0347 | 0.0452 | 0.0626 | 0.1047 | - | 0.0364 | 0.046 | 0.0653 | 0.1038 | - | ||||||||||||||||||
6.62% |
2.60% | |||||||||||||||||||
11.81% |
7.14% | |||||||||||||||||||
| Video & Games | Base | 0.0284 | 0.0353 | 0.0542 | 0.0816 | - | 0.0289 | 0.0355 | 0.0548 | 0.0810 | - | ||||||||||||||||||
8.59% |
5.36% | |||||||||||||||||||
6.81% |
7.00% | |||||||||||||||||||
| Software | Base | 0.0696 | 0.0895 | 0.1468 | 0.2264 | - | 0.0710 | 0.0893 | 0.1530 | 0.2258 | - | ||||||||||||||||||
2.16% |
5.23% | |||||||||||||||||||
4.06% |
4.68% | |||||||||||||||||||
| CDs & Vinyl | Base | 0.0148 | 0.0174 | 0.0317 | 0.0419 | - | 0.0149 | 0.0185 | 0.0326 | 0.0468 | - | ||||||||||||||||||
18.59% |
10.93% | |||||||||||||||||||
7.08% |
6.59% | |||||||||||||||||||
| Baby & Products | Base | 0.0112 | 0.0157 | 0.0260 | 0.0437 | - | 0.0109 | 0.0154 | 0.0257 | 0.0439 | - | ||||||||||||||||||
2.16% |
21.49% | |||||||||||||||||||
11.30% |
16.99% | |||||||||||||||||||
| Dataset | Method | UniSRec | MoRec | ||||||||||||||||||||||||||
| N@10 | N@20 | R@10 | R@20 | Avg. | N@10 | N@20 | R@10 | R@20 | Avg. | ||||||||||||||||||||
| Yelp | Base | 0.0380 | 0.0495 | 0.0737 | 0.1195 | - | 0.0391 | 0.0516 | 0.0757 | 0.1258 | - | ||||||||||||||||||
4.81% |
5.36% | |||||||||||||||||||
6.83% |
4.16% | |||||||||||||||||||
| Video & Games | Base | 0.0328 | 0.0421 | 0.0683 | 0.1054 | - | 0.0350 | 0.0438 | 0.0716 | 0.1065 | - | ||||||||||||||||||
3.97% |
7.76% | |||||||||||||||||||
0.08% |
2.64% | |||||||||||||||||||
| Software | Base | 0.0820 | 0.1041 | 0.1643 | 0.2522 | - | 0.0846 | 0.1050 | 0.1697 | 0.2510 | - | ||||||||||||||||||
2.49% |
1.30% | |||||||||||||||||||
2.96% |
4.37% | |||||||||||||||||||
| CDs & Vinyl | Base | 0.0150 | 0.0208 | 0.0298 | 0.0527 | - | 0.0186 | 0.0235 | 0.0405 | 0.0604 | - | ||||||||||||||||||
31.54% |
4.08% | |||||||||||||||||||
25.66% |
4.42% | |||||||||||||||||||
| Baby & Products | Base | 0.0152 | 0.0199 | 0.0315 | 0.0501 | - | 0.0176 | 0.0231 | 0.0371 | 0.0588 | - | ||||||||||||||||||
18.64% |
3.06% | |||||||||||||||||||
16.77% |
4.89% | |||||||||||||||||||
@@ 4.1.4. Implementation Details
We conduct all experiments on 8 NVIDIA A100 GPUs. To ensure a fair comparison, we set the embedding size and batch size for all methods to 256 and 2048, respectively. We optimize all models using the Adam (Kingma and Ba, 2014) optimizer with a learning rate of 0.001 and follow previous work (Sun et al., 2019) by adopting GeLU as the activation function. Following the existing works (Xie et al., 2022; Chen et al., 2022), we truncate user sequences to a maximum length of 50 across all datasets. Since our framework is model-agnostic, it can be seamlessly integrated into various sequential recommendation models. In particular, for BERT4Rec’s bidirectional Transformer, we employ a Prefix Masking strategy, where the item sequence part utilizes bidirectional attention, while the reasoning adopts unidirectional attention. Early stopping is triggered if the metrics on the validation set do not improve over 10 consecutive epochs. For item-based methods, we follow previous work (Liu et al., 2024) by using LLaMA-3.1-8B (Grattafiori et al., 2024) to encode item textual features. In particular, we apply Principle Component Analysis (PCA) to the averaged hidden states from the last layer, preserving core features and distilling 768-dimensional model representations. For ERL method, we search for the KL regularization hyperparameter within . For PRL method, we set the noise strength and tune the base temperature and temperature decay rate over the ranges and , respectively. Our codes will be available at https://github.com/TangJiakai/ReaRec.
@@ 5. Related Work
Sequential Recommendation. Sequential recommendation, as one of the core tasks within the field of Recommender Systems (RS), aims to predict the next item with which a user may interact by modeling behavior patterns and evolving user interests from user-item interaction sequences (Boka et al., 2024; Wang et al., 2019; Fang et al., 2020). Early studies focused on item-to-item collaborative transitions, typically employing Markov Chain-based matrix factorization methods (He and McAuley, 2016; Rendle et al., 2010) to achieve next-item predictions. With the advent of deep learning, researchers began exploring various sequential modeling architectures, such as Recurrent Neural Networks (RNNs) (Hidasi et al., 2015, 2016; Quadrana et al., 2017), Convolutional Neural Networks (CNNs) (Tang and Wang, 2018; Yuan et al., 2019; Yan et al., 2019), and Transformer (Kang and McAuley, 2018; Chen et al., 2019; Yang et al., 2022), to enhance modeling capabilities. Specifically, GRU4Rec (Hidasi et al., 2015) first introduced the GRU networks for session-based recommendation. Caser (Tang and Wang, 2018) applied convolutional operations by treating the item sequence embedding matrix as an “image” to extract multi-level interaction features. With the emergence of the Transformer architecture, numerous studies in sequential recommendation shifted toward self-attention-based models. For instance, SASRec (Kang and McAuley, 2018), a classic sequential recommendation baseline, incorporated self-attention to automatically learn the importance weights of historical items. BERT4Rec (Sun et al., 2019) further advanced this by adopting bidirectional encoding to capture more contextual dependencies from both directions of the sequence. Furthermore, to address data sparsity and cold-start issues, another line of research focused on leveraging common item attributes (such as text and images) to learn universal item and sequence representations (Hou et al., 2022; Yuan et al., 2023; Ding et al., 2021; Hou et al., 2023). For example, UniSRec (Hou et al., 2022) utilized multi-domain recommendation data to learn transferable sequential models. However, existing methods remain constrained by reasoning-free forward computation. In this paper, we further explore the potential reasoning capabilities of sequential recommenders by extending the multi-step computational depth at inference time.
@@
Inference-time Reasoning. As the scaling law of large language models (LLMs) at the training stage has gradually reached its bottleneck (Chen et al., 2025; Snell et al., 2024; Kumar et al., 2025), researchers have shifted their focus toward inference-time scaling. Notably, OpenAI’s O1 series (Jaech et al., 2024), Qwen’s QwQ series (Team, 2024), and DeepSeek’s R1 series (Guo et al., 2025) have emerged as key milestone works, marking the transition from conversational AI to reasoning-intensive AI and opening promising pathways toward achieving Artificial General Intelligence (AGI). These works leverage emerging long Chain-of-Thought (CoT) mechanisms to reveal excellent test-time scaling phenomena—where increased computational power (via generating more tokens) during inference substantially improves the model’s problem-solving abilities (Wei et al., 2022; Besta et al., 2024; Yao et al., 2023; Yang et al., 2024; Wang et al., 2022; Chu et al., 2023; Zhang et al., 2025). Extended reasoning space enables depth-scalable exploration, manifesting self-reflection capabilities (e.g., “Aha Moments”) that surpass traditional short-chain reasoning limitations (Chen et al., 2025; Ji et al., 2025). Compared with prior methods constrained by token-by-token generation in the discrete language spaces, which restricts the model’s expressive ability, another research direction focuses on implicit chain of thought reasoning in latent spaces, achieving both efficiency and performance gains in large language models (Hao et al., 2024; Xu et al., 2025a; Geiping et al., 2025; Chen et al., 2024) and multimodal foundation models (Shen et al., 2025; He et al., 2024). For example, Coconut (Hao et al., 2024) introduces continuous thinking in the latent reasoning space of LLMs, while Heima (Shen et al., 2025) compresses the entire multimodal CoT process into a single high-level thinking token (i.e., ¡CoT¿) to eliminate redundant intermediate token generation. Inspired by this think-before-action paradigm, we pioneer the exploration of implicit reasoning-enhanced sequential recommendation framework in this paper, proposing two lightweight reasoning learning methods to push the performance ceiling of sequential recommenders.